# Cracking Vector Search Indexes

### v1

@@ Abstract.
Retrieval Augmented Generation (RAG) uses vector databases to expand the expertise of an LLM model without having to retrain it. This idea can be applied over data lakes, leading to the notion of embeddings data lakes, i.e., a pool of vector databases ready to be used by RAGs. The key component in these systems is the indexes enabling Approximated Nearest Neighbor Search (ANNS). However, in data lakes, one cannot realistically expect to build indexes for every possible dataset. In this paper, we propose an adaptive, partition-based index, CrackIVF, that performs much better than up-front index building. CrackIVF starts answering queries by near brute force search and only expands as it sees enough queries. It does so by progressively adapting the index to the query workload. That way, queries can be answered right away without having to build a full index first. After seeing enough queries, CrackIVF will produce an index comparable to the best of those built using conventional techniques. As the experimental evaluation shows, CrackIVF can often answer more than 1 million queries before other approaches have even built the index and can start answering queries immediately, achieving 10-1000x faster initialization times. This makes it ideal when working with cold data or infrequently used data or as a way to bootstrap access to unseen datasets.
@@
PVLDB, 14(1): XXX-XXX, 2020.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 14, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
@@
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣setURL_TO_YOUR_ARTIFACTS.
@@ 1. Introduction
External data sources can complement Large language models (LLMs) (Brown et al., 2020) through the use of retrieval-augmented generation (RAG) (Lewis et al., 2020). In RAG, external data is represented as vector embeddings (i.e., learned representations of data that result in a multidimensional vector). The embeddings are indexed by a variety of means and searched through approximate nearest neighbor (ANN) search so that LLMs can answer queries on information they were not explicitly trained on. The effectiveness of RAG is tied to that of the ANN index structures, making them a critical component.
LLMs have been integrated with search engines (OpenAI, 2024) (e.g., Google, Bing, and Baidu). However, the vast repository of private, unstructured business data remains largely untapped. It has been estimated that unstructured data constitutes up to 80-90% of the total data volume (Merrill Lynch, 1998; MIT Sloan School of Management, 2021) and is often stored in open data formats on data lakes (Miller, 2018). However, as much as 70% of this data (Data Science Association, 2016) remains unused, classified as “dark” data (Heidorn, 2008; Zhang et al., 2016).
This data can be made discoverable (Chapman et al., 2020), but even then, the actual data often remains unindexed. For instance, as of March 2020, Google Research’s Dataset Search (Brickley et al., 2019) had indexed around 28 million structured and unstructured datasets (Benjelloun et al., 2020) based on their metadata. However, embedding-based vector search and RAG techniques make the promise to go a step further by enabling direct retrieval and question answering on the underlying data itself, inside of each dataset. This requires embedding the data, and the creation of approximate nearest neighbor (ANN) indexes for every single dataset, exposing the data to RAG systems. In this paper, we name such an approach embedding data lakes (EDL), where unstructured data is stored alongside its vector representations.
Efforts in this direction are already ongoing in both industry and academia. Databricks has deployed vector search capabilities (Databricks, 2024) over their Lakehouse (Armbrust et al., 2021), and new custom storage formats like Lance (LanceDB Developers, 2024) are being used to store embeddings on data lakes. Researchers are also directly exploring RAG-based techniques to query multi-modal data lakes (Tang et al., [n.d.]; Chen et al., 2023), as well as answering queries directly over unstructured data (Liu et al., 2025; Anderson et al., 2024). There is also work to directly define LLM-based data discovery systems over data lakes (An et al., 2025).
The issue we address in this paper is crucial to achieve the promise of embedding lakes (EDL): index selection at scale. EDL, in principle, would require building millions of ANN indexes across diverse datasets, modalities, workloads, and embeddings, making static index selection and pre-computation infeasible. In data lakes, some datasets may receive millions of queries, while others are rarely accessed. When selecting an index for embedding lakes, there is a trade-off. On the one hand, building an index requires significant upfront costs that might not pay off if the data is rarely used. On the other hand, skipping the indexing process and relying on a brute-force search is not scalable. Hence, data must be indexed using Approximate Nearest Neighbor (ANN) indexes but this raises two questions: when to index the data and how to index it.
@@
An optimal ANN index depends on workload patterns and future usage, which are often unknown. As shown in Figure 1, the index that minimizes total cumulative time spent on building and searching varies with the query volume. Larger indexes take longer to build and require many queries to justify their upfront cost (e.g., the large IVFlat configurations (5000, 16000 clusters) in the figure). Brute-force search (as shown in the figure), and smaller indexes (IVFlat 1000 clusters), minimize time-to-query for fast discovery but struggle at scale due to their higher query response times.
@@
To tackle index selection at scale, we adopt a well-established strategy: carefully avoiding the need to build the full index upfront. In this, we draw inspiration from deferred data structuring (Karp et al., 1988) and database cracking (Idreos et al., 2007; Halim et al., 2012; Schuhknecht et al., 2016), which have been successfully applied to single-dimensional database indexing. Similar approaches have been extended to multidimensional data (Holanda et al., 2018; Zardbani et al., 2023; Pavlovic et al., 2018; Nerone et al., 2021). However, prior work does not scale to the embedding dimensionalities nor supports the k-Nearest Neighbour (k-NN) queries typical in RAG systems. Recently, the AV-Tree was the first adaptive index for high-dimensional k-Nearest Neighbor (k-NN) search (Lampropoulos et al., 2023). AV-Tree targets short-lived data, with workloads of up to one thousand queries, and is a tree-based Exact Nearest Neighbor (ENN) adaptive index. This makes it impractical for RAG-scale workloads, which rely on Approximate Nearest Neighbor (ANN) index structures to trade off a bit of accuracy for large performance gains. Nonetheless, the idea of dynamically building an index tailored to the workload remains compelling. In this paper, we demonstrate how to achieve this efficiently in the context of RAG and employing the same data structures (Sivic and Zisserman, 2003) and systems (Douze et al., 2024) used for ANN search workloads (Aumüller et al., 2020; Simhadri et al., 2022) rather than analytics over cold data.
@@
We introduce CrackIVF, an incrementally built, partition-based ANN index that dynamically improves itself to adapt to increasing workload demands, minimizing both time-to-query and cumulative cost associated with suboptimal static index selection. CrackIVF evolves as a side effect of query execution. Each query is a candidate to add a new partition, or crack, while its search region is a candidate for local refinement, allowing the index to grow and improve as more queries arrive. Efficiency is maintained through two controls: one deciding where to apply cracking and refinement in the search space, and another controlling when, to balance indexing and search. In the paper we focus on a static setting (i.e., no data or query distribution shifts) and evaluate CrackIVF on multiple standard open source datasets for ANN search (Bertin-Mahieux et al., 2011; Pennington et al., 2014; Jégou et al., 2011; Babenko and Lempitsky, 2016).
@@
For partition-based indices, index selection means deciding upfront how many partitions the index will use, which has an impact on the ANN search performance. This decision has to be made without knowing the future query workload. CrackIVF waits to build the index until it sees enough queries and decides on the clusters based on the query distribution over the search space. As more queries arrive, the clusters are redefined, and their number is increased to adapt to the increasing query workload. Asymptotically, CrackIVF builds an index that converges to similar to those pre-built, but it has been able to answer queries along the way.
@@
Across benchmarks, CrackIVF consistently outperforms pre-built partition-based indexes. It converges to near-optimal query response time, yet achieves several orders of magnitude lower startup time than other indexes. It efficiently scales to large workloads and, in some cases, it can process 1 million queries before the baseline indexes have finished building. It is the only index that consistently remains near the Pareto frontier of minimum cumulative time across the number of queries posed to the system.
@@ 2. Background
In this section, we briefly cover the basics of approximated nearest neighbor search and IVF indexes.
Nearest Neighbor Search In the k-nearest neighbor (k-NN) search problem, we are given a set of points $P\in\mathbb{R}^{|P|\times d}$ and a query $q\in\mathbb{R}^{d}$. The aim is to find the $k$ points in $P$ that are closest to $q$ under some notion of distance or similarity (e.g., minimize the Euclidean distance (L2) or maximize the inner product (IP) to measure closeness). A straightforward Exact Nearest Neighbor (ENN) approach is a linear scan of every point, a.k.a brute-force search, incurring in $\mathcal{O}(|P|\cdot d)$ complexity. There exist more sophisticated exact indexing structures (e.g., KD-trees (Bentley, 1975), R-trees (Guttman, 1984), or VP-trees (Yianilos, 1993)). However, they often degrade to near-linear-scan performance in higher dimensions, due to the curse of dimensionality (Weber et al., 1998), making them impractical when $|P|$ or $d$ are large. With embeddings of hundreds to thousands of dimensions  (OpenAI, 2023) and real-world datasets at the scale of billions of data points (Simhadri et al., 2022), RAG systems turn to Approximate Nearest Neighbor (ANN) search techniques. They allow a small retrieval error $\varepsilon$ in exchange for better runtime or memory usage. This approximation quality can also be evaluated by the fraction of exact top-$k$ vectors recovered.
@@
Figure 2. Observations behind the design: (a) separating number of partitions and refinement of centroids; (b) access to the index is generally skewed towards certain regions; (c) regions queried often can be clustered and refined at a much lower granularity
Partition based IVF methods One of the most common ANN search structures is partition-based IVF indexes (Sivic and Zisserman, 2003; Chen et al., 2021; Sun et al., 2024). Instead of scanning all points in $P\subset\mathbb{R}^{d}$, they divide the vector space into $n_{\text{list}}$ disjoint groups, or partitions. At query time, the search process first identifies the $n_{\text{probe}}\leq n_{\text{list}}$ nearest partitions to $q$, whose representative vectors are closest (or most similar) to the query. Once these partitions are selected, the points within them are exhaustively scanned. Since the remaining partitions are skipped, some nearest neighbors may be missed, which is why the result is an approximate solution. A partition-based index is built by first selecting $n_{\text{list}}$ partition representative vectors, typically using $k$-means on a training subset $P_{\text{train}}\subset P$. Each data point is then assigned to the nearest representative, forming disjoint partitions. These assignments are typically stored in an inverted file (IVF) structure (Sivic and Zisserman, 2003). A widely adopted implementation of IVF indexes is the FAISS library from Meta (Douze et al., 2024), which we use as a baseline. See Section 6 for other ANN index types.
@@ 3. Solution Overview
We focus on Partition-based IVF indexes because they offer several advantages for deployment in embedding data lakes. They offer the best trade-off in index build time, query response time, and index size (Research, 2023). They have the lowest index build times relative to query throughput (QPS). All of the above are crucial for embedding data lakes, where potentially millions of indexes of various sizes and levels of utilization will have to be constructed and stored.
@@ 3.1. Cracking in high dimensions
A well known approach to avoid upfront cost for indexing or organizing data is cracking. In database cracking (Idreos et al., 2007), there is a final state where a single column is sorted. Over time, fewer points are moved, and over shorter distances, reducing the overhead of the cracking. Additionally, in a naive implementation, the physical reorganization happens with every incoming query.
In high-dimensional spaces, the assumption of a fully sorted state no longer holds, and space filling curves, like Z-order (Samet, 1984), do not scale past a few dimensions. Additionally, enforcing reorganization after every query becomes increasingly difficult, as it is a more expensive operation. Moving points between inverted lists incurs random access overhead, with no eventual locality benefits. High-dimensional vectors require moving multiple bytes per dimension, increasing data movement costs. Furthermore, inverted lists, are often implemented as memory-aligned contiguous arrays to allow for efficient SIMD-based kernels, and they may require linear-time compaction even for a single vector to move. These factors contribute to the significant latency overhead, making physical reorganization after every query difficult.
Based on this, the challenge to address is as follows A cracking-based IVF index must efficiently manage the significantly higher overhead of physical reorganization in high dimensions, while minimizing time-to-query, continuously improving the index, and ensuring minimum cumulative time across all scales.
@@ 3.2. Outline of the approach
As Figure 2(a) illustrates, index construction can be divided into two operations that independently improve index performance: (i) increasing the number of partitions by randomly selecting partition representatives and (ii) refining these representatives into well-placed centroids using k-means training, which helps to distribute points more evenly across partitions and improve the overall representativeness. Although both improve the index independently, the best results consistently come after reaching a maximum index size, even with randomly initialized representatives, followed by k-means training to optimize their placement. Thus, the optimal approach is to first increase the number of partitions up to a limit and then refine their locations using k-means.
@@
Observation 1 behind the approach: Index construction can be decoupled into two distinct build operations, which can improve the index independently and be separated in time. .
@@
Existing partition-based IVF vector search indexes apply build operations evenly across the vector space, with centroids sampled uniformly and K-means refinements performed simultaneously throughout. While simple and intuitive, this approach assumes equal importance across the space, and requires that both operations are applied at the same time on the entire space leading to larger than necessary index build times. However, vector search workloads can exhibit highly skewed access patterns, where certain regions are significantly more important than others. For instance, in an industrial workload (Mohoney et al., 2023, 2024), up to 85% of centroids are never accessed during search, which would make partitioning or training efforts in those regions unnecessary. As we show in Figure 2(b), skewed query distributions also appear in well known open-source vector search datasets (Bertin-Mahieux et al., 2011; Pennington et al., 2014; Jégou et al., 2011; Babenko and Lempitsky, 2016). Even on open source datasets showcasing moderate amounts of skew, where $80\%$ of the total queries access $60\%$ of the partitions, the disparity between the most frequently and least frequently accessed partitions is upwards of 50 times. There is a fundamental inefficiency in current uniform indexing approaches: a large portion of the vector space may rarely accessed, yet the indexing effort is evenly distributed across the space, and always paid upfront.
@@
Observation 2 behind the approach: Just as index construction can be separated in time, it can also be separated in space. The index build operations do not have to be applied uniformly across the vector space; instead, they can be focused on regions that are accessed more frequently. Since certain areas of the vector space are more critical than others, adapting index construction to follow the query distribution is a natural strategy to optimize efficiency and accelerate indexing.
Lastly, in partition-based IVF indexes, a query first identifies the nearest partitions and then scans all points within those partitions to retrieve the $k$ nearest neighbors. This involves the same distance computations needed to update the index, which naturally leads to a cracking-based approach, where index refinement occurs as a side effect of query execution.
@@
Observation 3 behind the approach: We can further amortize the cost of specific build operations, such as assigning vectors to partitions and identifying the local regions to crack and refine, Figure 2(c). We can reuse computations from search operations to perform index construction as a side effect of query execution.

### v2

@@ Abstract.
Retrieval Augmented Generation (RAG) uses vector databases to expand the expertise of an LLM model without having to retrain it. The idea can be applied over data lakes, leading to the notion of embedding data lakes, i.e., a pool of vector databases ready to be used by RAGs. The key component in these systems is the indexes enabling Approximated Nearest Neighbor Search (ANNS). However, in data lakes, one cannot realistically expect to build indexes for every dataset. Thus, we propose an adaptive, partition-based index, CrackIVF, that performs much better than up-front index building. CrackIVF starts answering as a small index, and only expands to improve performance as it sees enough queries. It does so by progressively adapting the index to the query workload. That way, queries can be answered right away without having to build a full index first. After seeing enough queries, CrackIVF will produce an index comparable to those built with conventional techniques. CrackIVF can often answer more than 1 million queries before other approaches have even built the index, achieving 10-1000x faster initialization times. This makes it ideal for cold or infrequently used data and as a way to bootstrap access to unseen datasets.
@@
PVLDB, 18(11): 3951 - 3964, 2025.
doi:10.14778/3749646.3749666 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 11 ISSN 2150-8097.
doi:10.14778/3749646.3749666
@@
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/mageirakos/crack-ivf-vldb.
@@ 1. Introduction
Large language models (LLMs) (Brown et al., 2020) can be complemented with retrieval-augmented generation (RAG) (Lewis et al., 2020). In RAG, external data is represented as vector embeddings (i.e., learned representations of data that results in a multidimensional vector). It is indexed and accessed through approximate nearest neighbor (ANN) search so that LLMs can answer queries on information they were not trained on. The effectiveness of RAG is tied to that of the ANN index structures, making them a critical component.
@@
In practice, the vast majority of private, unstructured data remains untapped. It is estimated to be up to 80-90% of all data (Merrill Lynch, 1998; Harbert, 2021), and is often stored in data lakes on open data formats (Miller, 2018). As much as 70% of it remains unused (Gualtieri, 2016), classified as “dark” data (Heidorn, 2008; Zhang et al., 2016). While the datasets can be made discoverable (Chapman et al., 2020), the actual underlying data in them often remains unindexed. As of March 2020, Google Research’s Dataset Search (Brickley et al., 2019) had indexed around 28 million structured and unstructured datasets (Benjelloun et al., 2020) based on their metadata. Embedding-based vector search and RAG techniques could however help us go a step further by enabling direct retrieval and question answering on the underlying data itself. This requires embedding the data, as well as the creation of approximate nearest neighbor indexes so as to expose the data to RAG systems. We refer to such an approach as embedding data lakes (EDL), where unstructured data is stored alongside its vector representations and is queried by a pool of vector databases. Efforts in this direction are already ongoing. Databricks uses vector search (Databricks, 2024) over their Lakehouse (Armbrust et al., 2021), and new custom storage formats (LanceDB Developers, 2024) are being used to store embeddings on data lakes. Researchers are also exploring RAG-based techniques to query multi-modal data lakes (Tang et al., 2024; Chen et al., 2023), as well as answering queries over unstructured data (Liu et al., 2025; Anderson et al., 2024).
The issue we address in this paper is index selection at scale. Embedding data lakes, in principle, would require building millions of indexes across diverse datasets, modalities, workloads, and embedding models, which is neither feasible nor cost efficient. Some datasets may receive the majority of queries while others are rarely accessed. When selecting an index for embedding lakes, there is a trade-off. On the one hand, building an index requires significant upfront costs that might not pay off if the data is rarely queried. On the other hand, skipping the indexing process and relying on a brute-force search is not scalable. Hence, data must be indexed, which raises two questions: when and how to index the data.
@@
An optimal ANN index depends on the workload patterns and future usage, which are often unknown. As shown in Figure 1, the index that minimizes total cumulative time spent on building and searching varies with the query volume. Larger indexes take longer to build and require many queries to justify their upfront cost (e.g., the large IVFlat configurations (5000, 16000 clusters) in the figure). Brute-force search, and smaller indexes (IVFlat 1000 clusters), minimize time-to-query for fast discovery but struggle at scale due to their higher query response times.
@@
To tackle index selection at scale, we adopt a well-established strategy: carefully avoiding the need to build the full index upfront. We draw inspiration from deferred data structuring (Karp et al., 1988) and database cracking (Idreos et al., 2007; Halim et al., 2012; Schuhknecht et al., 2016). Similar approaches have been extended to multidimensional data (Holanda et al., 2018; Zardbani et al., 2023; Pavlovic et al., 2018; Nerone et al., 2021). However, prior work does not scale to the embedding dimensionalities nor supports the k-Nearest Neighbour (k-NN) queries typical in RAG deployments. Recently, the AV-Tree was the first adaptive index for high-dimensional k-NN search (Lampropoulos et al., 2023). AV-Tree targets short-lived data, with workloads of up to one thousand queries, and is a tree-based Exact Nearest Neighbor (ENN) adaptive index. This makes it impractical for RAG-scale workloads, which rely on Approximate Nearest Neighbor index structures to trade off a bit of accuracy for lower query response time and higher scalability. Nonetheless, the idea of dynamically building an index tailored to the workload is compelling. In this paper, we demonstrate how to do this efficiently in the context of RAGs by employing the same data structures (Sivic and Zisserman, 2003) and systems (Douze et al., 2024) used for ANN search workloads (Aumüller et al., 2020; Simhadri et al., 2022).
@@
We introduce CrackIVF, an incrementally built, partition-based ANN index that dynamically improves itself to adapt to increasing workload demands, minimizing both time-to-query and cumulative cost associated with suboptimal static index selection. CrackIVF evolves as a side effect of query execution. Each query is a candidate around which we can add a new partition, or crack. Likewise, its local region visited during search is a candidate for local refinement, allowing the index to grow and improve as more queries arrive. Efficiency is maintained through two control mechanisms: one deciding where to apply cracking and refinement in the search space, and another controlling when, balancing indexing and search times. In this work we focus on the static setting (i.e., no data or query distribution shifts) and evaluate CrackIVF on multiple standard open source datasets for ANN search (Bertin-Mahieux et al., 2011; Pennington et al., 2014; Jégou et al., 2011; Babenko and Lempitsky, 2016).
@@
For partition-based indices, index selection means deciding upfront how many partitions the index will use, which has an impact on the ANN search performance. More partitions increase the upfront construction cost but improve query performance. This decision has to be made without knowing the future query workload. On the other hand, CrackIVF waits to build the index as it sees enough queries and decides on the clusters based on the query distribution over the search space. As more queries arrive, the clusters are redefined, and their number is increased to adapt to the increasing query workload. CrackIVF is an incrementally built index that asymptotically matches or surpasses the performance of pre-built indexes, while it has been able to answer queries along the way.
@@
Across benchmarks, CrackIVF consistently outperforms pre-built partition-based indexes. It converges to near-optimal query response time, yet achieves several orders of magnitude lower startup time than other indexes. It efficiently scales to large workloads and, in some cases, can process 1 million queries before the baseline indexes have finished building. It is the only index that consistently remains near the Pareto frontier of minimum cumulative time across the number of queries posed to the system.
@@ 2. Background
Nearest Neighbor Search In the k-nearest neighbor (k-NN) search problem, we are given a set of points $P\in\mathbb{R}^{|P|\times d}$ and a query $q\in\mathbb{R}^{d}$. The aim is to find the $k$ points in $P$ that are closest to $q$ under some notion of distance or similarity (e.g., minimize the Euclidean distance (L2) or maximize the inner product (IP)). A straightforward Exact Nearest Neighbor approach is a linear scan of every point, a.k.a brute-force search, incurring $O(|P|\cdot d)$ complexity. There exist more sophisticated exact indexing structures (e.g., KD-trees (Bentley, 1975), R-trees (Guttman, 1984)). However, due to the curse of dimensionality (Weber et al., 1998), performance degrades to near-linear-scan in higher dimensions, making them impractical when $|P|$ or $d$ are large. With embeddings of hundreds to thousands of dimensions  (OpenAI, 2025) and real-world datasets at the scale of billions of data points (Simhadri et al., 2022), RAG systems turn to Approximate Nearest Neighbor search techniques. They allow a small retrieval error $\varepsilon$ in exchange for better runtime or memory usage. The approximation quality can be evaluated by the fraction of exact top-$k$ vectors recovered.
@@
Figure 2. Observations behind the design: (a) separating the number and refinement of partitions; (b) access to the index is generally skewed towards certain regions; (c) regions queried often can be clustered and refined at a much lower granularity
Partition based IVF methods One of the most common ANN search structures is partition-based IVF indexes (Sivic and Zisserman, 2003; Chen et al., 2021; Sun et al., 2024). Instead of scanning all points in $P\subset\mathbb{R}^{d}$, they divide the vector space into $n_{\text{list}}$ disjoint groups, or partitions. At query time, the search process first identifies the $n_{\text{probe}}\leq n_{\text{list}}$ nearest partitions to $q$, whose representative vectors are closest (or most similar) to the query. Once these partitions are selected, the points within them are exhaustively scanned. Since the remaining partitions are skipped, some nearest neighbors may be missed, which is why the result is an approximate solution. A partition-based index is built by first selecting $n_{\text{list}}$ representative vectors for each partition, typically using $k$-means on a training subset $P_{\text{train}}\subset P$. Each data point is then assigned to the nearest representative. These assignments are typically stored in an inverted file (IVF) structure (Sivic and Zisserman, 2003). A widely adopted implementation of IVF indexes is the FAISS library from Meta (Douze et al., 2024), which we built CrackIVF on and use as a baseline. See Section 6 for other ANN index types.
@@ 3. Solution Overview
[…]
@@ 3.1. Cracking in high dimensions
An approach to avoid the upfront cost for indexing is cracking. In the original database cracking (Idreos et al., 2007), the final state is where a single column is sorted. Over time, fewer points are physically reorganized reducing the overhead. In high-dimensional spaces, the having a final sorted state no longer holds, and space filling curves, like Z-order (Samet, 1984), do not scale past a few dimensions. Additionally, in a naive implementation, the physical reorganization happens with every incoming query. In high-dimensions, this becomes increasingly difficult. High-dimensional vectors require moving multiple bytes per dimension, increasing data movement costs. In IVF, moving points between inverted lists incurs random access overhead, with no eventual locality benefits. Furthermore, inverted lists are often implemented as memory-aligned contiguous arrays to allow for efficient SIMD-based operations, and they may require linear-time compaction even for a single vector to move. These factors contribute to the latency overhead, making physical reorganization after every query difficult. The challenge is as follows: A cracking-based IVF index must efficiently manage the significantly higher overhead of physical reorganization in high dimensions, while minimizing time-to-query, continuously improving the index, and ensuring minimum cumulative time across all scales.
@@ 3.2. Outline of the approach
Index construction can be divided into two operations that independently improve index performance, i) increasing the number of partitions and ii) refining partitions by using K-means, which helps to distribute points more evenly (Figure 2(a)). The best results consistently come after reaching a maximum index size, even with randomly chosen representatives, followed by K-means training to optimize their placement. Thus, the optimal approach is to first increase the number of partitions up to a limit and then also refine their locations using K-means.
@@
Observation 1: Index construction can be decoupled into two distinct build operations, which can improve the index independently and be separated in time.
@@
Existing partition-based IVF vector search indexes apply build operations evenly across the vector space, with partitions sampled uniformly and K-means refinements performed globally. While simple and intuitive, this approach assumes equal importance across the space, leading to larger than necessary index build times. However, vector search workloads can exhibit highly skewed access patterns, where certain regions are significantly more important than others. For instance, in an industrial workload (Mohoney et al., 2023, 2024), up to 85% of partitions are never accessed during search, which would make partitioning or training efforts in those regions unnecessary. As we show in Figure 2(b), skewed query distributions also appear in well known open-source vector search datasets (Bertin-Mahieux et al., 2011; Pennington et al., 2014; Jégou et al., 2011; Babenko and Lempitsky, 2016). Even on open source datasets showcasing moderate amounts of skew, where $80\%$ of the total queries access $60\%$ of the partitions, the disparity between the most frequently and least frequently accessed partitions is upwards of 50 times. There is a fundamental inefficiency in current uniform indexing approaches. A large portion of the vector space may rarely accessed, yet the indexing effort is evenly distributed across the space, and always paid upfront.
@@
Observation 2: Just as index construction can be separated in time, it can also be separated in space. Index build operations do not have to be applied uniformly across the vector space. Instead, they can be local to regions that are accessed. Adaptive index construction that follows the query distribution is one strategy to optimize for efficiency.
In partition-based IVF indexes, operations performed during search and those performed during index construction are very similar. During search the distances of the query to centroids and nearby points are computed, and during index construction the distances of points to centroids are computed. This means that by treating the query as a potential new partition representative “centroid”, we can both identify good locations to add new cracks, and reuse the intermediate computations of search to perform incremental index construction. This naturally leads to a cracking-based approach for IVF, where index construction can occur as a side effect of query execution.
@@
Observation 3: We can further amortize the cost of specific build operations, such as assigning vectors to partitions and identifying the local regions to crack and refine, Figure 2(c). We can reuse computations from search operations to perform index construction as a side effect of query execution.